OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at the origin and a radius of 9? (2 points) x^2 + (y - 9)^2 = 1 (x - 9)^2 + y^2 = 1 x^2 + y^2 = 9 x^2 + y^2 = 81

OpenStudy (anonymous):

for x^2 + (y - 9)^2 = 1 (0,9) , r =1 for (x - 9)^2 + y^2 = 1 (9,0) , r =1 for x^2 + y^2 = 9 (0,0) , r =3 for x^2 + y^2 = 81 (0,0) , r = 9

OpenStudy (anonymous):

the general equation of any circle is (x-a)^2+(y-b)^2=r^2 where (a,b)=the center coordinates r=the radius

OpenStudy (anonymous):

Ohh Okay , I Understand , Thanks Guys @esamalaa @hazem-wardat